A medical technician is working with the four samples of radionuclides listed in the table below. Initially, each sample contains 14.00 ?mol of the radionuclide First, order the samples by decreasing initial radioactivity. Then calculate how long it will take for the amount of radionuclide in each sample to decrease to 1/16 of the initial amount. radionuclide initial radioactivity time for amount of radionuclide to decrease to 1/16 of initial amount sample symbol half-life days days days 25.0 days (choose one) 15 1 (highest) 91 39 59.0 days 4 (lowest) 84.0 days (choose one) 21 122 53 4.0 minutes (choose one) minutes
Answer #1

Answer -

Initial radioactivity of those elements are hightest which have small value of half life.

Hence order of initial radioactivity -D(highest)>A>B>C(lowest)

Part B -

Here initial Radioactivity problem is right .

Now we solve for time for amount of radionuclide to decrease to 1/16 of iniyial amount.

After 4 half life it is 1/16 of initial amount.

A. It takes 4 hlaf life to 1/16 of initial amount .

hence , time = 4*25 =100 days

B. It takes 4 hlaf life to 1/16 of initial amount .

hence , time = 4*59 = 236 days

C. It takes 4 hlaf life to 1/16 of initial amount .

hence , time = 4*84 = 336 days

D. It takes 4 hlaf life to 1/16 of initial amount .

hence , time = 4*4 =16 days